# Careful finite element simulations of cold rolling with accurate through-thickness resolution and prediction of residual stress

### v1

@@ 6 Conclusion
We show that our through-thickness predictions compare well with a rigid-perfectly-plastic analytical model [13]. When optimized correctly, analytical mathematical models can be used in online control due to their rapid computation time and ability to provide accurate results compared to FE simulations. Excellent agreement in shear stress was found when comparing the analytical and FE models. Both models captured the oscillatory nature of shear during rolling. The relatively small discrepancies between the shear results of the two models can be attributed to the size of the roll-gap aspect ratio, $\varepsilon$ and the differences in the material and friction models. In the limits as $\varepsilon\rightarrow 0$, $\gamma\rightarrow 0$ and $E\rightarrow\infty$, we expect that the agreement between the FE and analytical results would be even further improved.
@@
Without the ability to compare to an analytical model, previous studies have often relied on the convergence of roll force and roll torque to assess the quality of numerical rolling simulations. Here, we show that roll force and roll torque are poor indicators of simulation quality. This is because these averaged quantities often converge well (see Figure 4(b)) even for under-resolved simulations that fail to accurately predict other quantities of interest (such as the distribution of shear stress at the roll exit in Figure 5(b)). Comparison to a steady state analytical model does however require that time-domain FE simulations converge to a steady state. Here, instead of starting our simulations with the sheet in front of the rollers and pushing the sheet into the roll gap, as is common practice, we start with the sheet in the roll gap but with the rollers removed and then squeezed the rollers into the sheet in a preliminary bite step before rolling (as described in subsection 3.3). This was found to improve convergence to a steady state of rolling.
@@
Unexpectedly, our results show the existence of a region near the neutral point where the sheet and rollers are sticking, rather than slipping. This is not commonly described in previous studies, even in studies of extreme cases such as foil rolling where many different regions within the roll gap are detailed and analysed (e.g., Fleck et al. [64], Figure 3c, zones E and F). The assumption that only slipping occurs at the neutral point is a common one, and indeed it is the cause of the discrepancy between the present FE model and the analytical model. While care has been taken here to ensure that the sticking region near the neutral point is not a numerical artefact, it is a consequence of Coulomb friction, which may be an overly-simplified model of the friction that practically occurs in metal rolling. Comparison with the analytical model shows that the effects of this sticking region on the stress oscillation pattern through-thickness are important beyond the neutral point and affects the shear stresses at the roll-gap exit.
@@ Acknowledgements
AOC is supported by the European Union, Science Foundation of Ireland and Lero, the Science Foundation Ireland Research Centre for Software, grants #101028291 #13/RC/2094 and #SOWP2-TP0023, respectively. ME gratefully acknowledges the funding of a University of Warwick Chancellor’s Scholarship. EJB is grateful for the UKRI Future Leaders’ Fellowship funding (MR/V02261X/1) supporting this work. DOK is grateful for funding from Science Foundation Ireland (21/FFP-P/10160).
Supplementary Material
Appendix S1 ABAQUS formulation
@@
S1.1 ABAQUS/Standard formulation
ABAQUS/Standard is an implicit solver where the model is updated from time $t$ to time $t+\Delta t$ using information from times $t$ and $t+\Delta t$. In this work, only static analyses are considered in ABAQUS/Standard. Therefore, since inertia is ignored, the equations of motion are now
@@
${\Delta\mathbf{u}_{\left(n+1\right)}}^{\left(t+\Delta t\right)}=\left[{\mathbf{K}_{\left(n\right)}}^{\left(t+\Delta t\right)}\right]^{-1}\cdot\left({\mathbf{F}_{\left(n\right)}}^{\left(t+\Delta t\right)}\right),$ (3)
${\mathbf{u}_{\left(n+1\right)}}^{\left(t+\Delta t\right)}={{\mathbf{u}}_{(n)}}^{\left(t+\Delta t\right)}+\Delta{\mathbf{u}_{\left(n+1\right)}}^{\left(t+\Delta t\right)},$ (4)
where the superscript $\left(t+\Delta t\right)$ is the time at which the equations are evaluated, the subscripts $\left(n\right)$ and $\left(n+1\right)$ indicate the current and next iterations respectively and $\mathbf{K}_{\left(n\right)}=\left(\partial\mathbf{F}/\partial\mathbf{u}\right)_{\left(n\right)}$ is the stiffness matrix at iteration $n$. The inversion of the stiffness matrix $\mathbf{K}_{\left(n\right)}$ is computationally expensive. The convergence of the Newton–Raphson method is measured by ensuring that all entries $\mathbf{F}(\mathbf{u})$ and $\Delta\mathbf{u}$ are sufficiently small [66]. At the end of time increment $t$, the next time increment $t+\Delta t$ is attempted, and load or displacement is incremented corresponding to $\Delta t$. If convergence is not achieved by ABAQUS in a certain time increment, $\Delta t$ is reduced and the Newton-Raphson method is attempted again. The reader should note that no velocities or accelerations are given by ABAQUS/Standard for static analyses. These can be manually extracted via postprocessing of the displacements (see Section S2).
S1.2 ABAQUS/Explicit formulation
@@
$\dot{\mathbf{u}}^{\left(i+\frac{1}{2}\right)}=\dot{\mathbf{u}}^{\left(i-\frac{1}{2}\right)}+\frac{\Delta t^{\left(i+1\right)}+\Delta t^{\left(i\right)}}{2}\ddot{\mathbf{u}}^{(i)},$ (6)
${\mathbf{u}}^{\left(i+1\right)}={\mathbf{u}}^{\left(i\right)}+\Delta t^{\left(i+1\right)}\dot{\mathbf{u}}^{\left(i+\frac{1}{2}\right)}.$ (7)
@@
where $L^{e}$ is the smallest element dimension in the FE model, $\rho$ is the material density and $E$ is the Young’s modulus of the material. However, ABAQUS/Explicit enables the use of mass scaling. Mass scaling uniformly scales the density of the material within the simulation to facilitate longer time steps and reduce computational time [43]. The limited time increment means that explicit methods are conditionally stable, thus the mass scaling factor (MSF) must be chosen carefully to reduce simulation times without introducing inertial effects. The introduction of inertial effects can lead to erroneous and unstable outputs [68, 69, 57]. The general rule of thumb to assess inertial effects is to ensure that the ratio between the kinetic energy and the internal energy is small (5-10%) [67, 15].
Appendix S2 FE velocity calculation
ABAQUS/Standard does not output velocity components or PEEQ rate values for static analyses. Therefore, we manually calculate the velocity and PEEQ rate values used to produce Figure 9. The velocity and PEEQ rate approximations are conducted via a Lagrangian time derivative of the displacements and the PEEQ values respectively. Since the procedure is identical for calculating both velocity and PEEQ rate values, we give details of the velocity calculation only.
@@
${v}^{(f)}_{x}=\frac{\left(\frac{1}{12}\right){u}^{(f-2)}_{x}-\left(\frac{2}{3}\right){u}^{(f-1)}_{x}+\left(\frac{2}{3}\right){u}^{(f+1)}_{x}-\left(\frac{1}{12}\right){u}^{(f+2)}_{x}}{\Delta t},$ (9a)
${v}^{(f)}_{z}=\frac{\left(\frac{1}{12}\right){u}^{(f-2)}_{z}-\left(\frac{2}{3}\right){u}^{(f-1)}_{z}+\left(\frac{2}{3}\right){u}^{(f+1)}_{z}-\left(\frac{1}{12}\right){u}^{(f+2)}_{z}}{\Delta t},$ (9b)
@@
where $\Delta t$ is constant between all time frames considered. In order to allow velocities and all other stress and strain components to be evaluated at the same time frame, all data in Section 5 is evaluated at a time point that is two time frames before $t=0.1\,$\mathrm{s}$$, which is the end of the rolling step. The authors ensured that $\Delta t$ is sufficiently small for accurate velocity approximations.
Appendix S3 Incompatible CPE4I elements
@@
Figure S1 compares interpolated shear stress from two simulations: one with CPE4I elements and the other with CPE4R elements. Both simulations employ $N_{e}=15$ elements through the half-thickness, since simulations with $N_{e}\geq 20$ CPE4I elements required extremely small time increments which gave rise to convergence issues. The maximum relative error in Figure S1 is approximately 9% when compared to the maximum absolute shear stress in the roll gap. We expect that this error would reduce significantly as the mesh density increases. Given the expected similar shear stresses with increasing mesh density and that the simulation with CPE4I elements requires 17.7 CPU hours compared to 3.2 CPU hours for the simulation with CPE4R elements, CPE4I elements were not used for the main results in this work.
@@
Figure S1: Interpolated results for shear stress ($\mathrm{M}\mathrm{P}\mathrm{a}$) from simulations with CPE4I elements and CPE4R elements. Both simulations employ $N_{e}=15$ elements, a radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$. The results are almost identical and the CPU time is 17.7 hours for the CPE4I simulation compared to 3.2 hours for the CPE4R simulation. The solid, dotted and dashed lines depict the interpolated results in the sheet at $x/L=0.025$, $x/L=0.6913$ and $x/L=1.09$ respectively (see Figure 1). The horizontal centre (symmetry axis in Figure 1) of the sheet corresponds to $z/h_{0}=0$.
Appendix S4 Relative error: choice of $x/L$ locations
@@
We detail here how the three $x/L$ positions used in the interpolated results throughout this paper were chosen. These $x/L$ positions represent the positions of maximum relative errors. To explain how these errors are calculated, let us consider the solid-line shear results in Figure 5b. The interpolated shear result from the simulation with $N_{e}=5$ is interpolated onto a curve with 40 points. The absolute difference between this new curve and the interpolated shear result from the simulation with $N_{e}=40$ is calculated at each $z/h_{0}$ point. Finally, the absolute difference at each point is divided by the maximum absolute shear stress in the entire roll gap, as predicted by the simulation with $N_{e}=40$. At each $x/L$ location, the maximum relative error value out of the 40 values is stored in an array. The same procedure is carried out for $N_{e}=10$, $N_{e}=15$, $N_{e}=20$ and $N_{e}=30$, and all of these maximum errors are plotted in Figure S2.
@@
Figure S2: Relative error in (a) von Mises stress, (b) shear stress and (c) PEEQ (plastic equivalent strain) from simulations with varying mesh densities. All simulations are compared to a simulation with $N_{e}=40$ elements through the half-thickness. The error value at each $x/L$ location represents the maximum error through the thickness of the sheet at that $x/L$ position. All simulations employ a radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
@@
Figure S2 shows that $x/L=0.025$ is a position where high relative errors occur, and hence this is one of the $x/L$ positions used for interpolations in this paper. Figure 2(a) and Figure 2(b) also show a spike in relative errors near $x/L=1.09$. This is another $x/L$ position used to calculate the interpolated results. It is not clear from Figure 2(b) and Figure 2(c) which third $x/L$ position to choose. However, Figure 2(a) shows a small peak in relative error around the neutral point $x/L=0.6913$ (as predicted by the implicit simulation with $N_{e}=40$). Given this peak and the interesting dynamics around the neutral point in general, $x/L=0.6913$ is the third position used in the interpolated results throughout this paper.
Appendix S5 Slip line theory
@@
$\phantom{=}\quad\quad\quad\,=\begin{pmatrix}S_{xx}-p&S_{xz}\\ S_{xz}&S_{zz}-p\end{pmatrix},$
@@
$0=\frac{\partial\sigma_{xx}}{\partial x}+\frac{\partial\sigma_{xz}}{\partial z}=-\frac{\partial p}{\partial x}-\frac{2Y}{\sqrt{3}}\left(\cos{2\theta}\frac{\partial\theta}{\partial x}+\sin{2\theta}\frac{\partial\theta}{\partial z}\right),$ (14)
$0=\frac{\partial\sigma_{xz}}{\partial x}+\frac{\partial\sigma_{zz}}{\partial z}=-\frac{\partial p}{\partial z}-\frac{2Y}{\sqrt{3}}\left(\sin{2\theta}\frac{\partial\theta}{\partial x}-\cos{2\theta}\frac{\partial\theta}{\partial z}\right).$ (15)
Equations 14 and 15 are hyperbolic PDEs and can be solved by the method of characteristics, (see Johnson et al. [47] for a detailed description). The characteristic directions of each of these equations are given by
$\dfrac{dz}{dx}=\tan{\theta},$ (16)
$\dfrac{dz}{dx}=\tan{\left(\theta+\frac{\pi}{2}\right)},$ (17)

### v2

@@ 7 Conclusion
We show that our through-thickness predictions compare well with a rigid-perfectly-plastic analytical model [21]. Excellent agreement in shear stress was found when comparing the analytical and FE models. Both models captured the oscillatory nature of shear during rolling. The relatively small discrepancies between the shear results of the two models can be attributed to the size of the roll-gap aspect ratio, $1/\varepsilon$ and the differences in the material and friction models, with the analytic model not allowing for a sticking region near the neutral point. In the limits as $\varepsilon\rightarrow 0$, $\gamma\rightarrow 0$ and $E\rightarrow\infty$, we hypothesise that the agreement between the FE and analytical results would be even further improved.
@@
Without the ability to compare to an analytical model, previous studies have often relied on the convergence of roll force and roll torque to assess the quality of numerical rolling simulations. Here, we show that roll force and roll torque are poor indicators of simulation quality. This is because these averaged quantities often converge well even for under-resolved simulations (see $N_{e}=5$ results in Figure 6) that fail to accurately predict other quantities of interest such as the through-thickness roll-gap shear distribution (see $N_{e}=5$ results in Figure 7(b)). Comparison to a steady-state analytical model does however require that time-domain FE simulations converge to a steady state. Here, instead of starting our simulations with the sheet in front of the rollers and pushing the sheet into the roll gap, as is common practice, we started with the sheet in the roll gap but with the rollers removed and then squeezed the rollers into the sheet in a preliminary bite step before rolling (as described in subsection 3.3). This was found to improve convergence to a steady state of rolling. We confirm that a steady state has been reached via stagnant through-thickness stress and strain quantities, not via roll force and torque, as conducted by previous authors. Although the roll force and torque are necessarily constant during the steady state of the rolling process, constant roll force and torque are not sufficient for identifying steady-state behaviour on their own, as shown here.
@@
Unexpectedly, our results show the existence of a region near the neutral point where the sheet and rollers are sticking, rather than slipping. This is not commonly described in previous studies, even in studies of extreme cases such as foil rolling where many different regions within the roll gap are detailed and analysed (e.g., Fleck et al. [71], Figure 3c, zones E and F). The assumption that only slipping occurs at the neutral point is a common one, and indeed this assumption is the cause of the analytic model deviating from the FE model near the netural point. While care has been taken here to ensure that the sticking region near the neutral point is not a numerical artefact, it is a consequence of Coulomb friction, which may be an overly-simplified model of the friction that practically occurs in metal rolling. Comparison with the analytical model shows that the effects of this sticking region on the stress oscillation pattern through-thickness are important beyond the neutral point and affects the shear stresses at the roll-gap exit.
@@
All of the analysis here, and many of the previous studies in Table 1, assume plane strain to reduce the problem to 2D. This is valid sufficiently far from the edge of the sheet provided the sheet is sufficiently wide and the rolls are perfectly aligned. However, in order to study edge effects, misalignment of the rolls, or less-wide sheets, a fully 3D study would be needed. A consequence of the analysis here is that $602N_{e}=$ 60 elements through-thickness were needed to accurately resolve through-thickness behaviour, and this level of detail would be extremely computationally-expensive in a 3D simulation, especially since, to avoid elements with a large aspect ratio, a comparably large number of across-width elements would presumably also be needed. It may be possible for very narrow strips to instead assume a plane stress condition and again reduce the problem to a (different) 2D model, although this has not been investigated further here.
@@ Acknowledgements
ANOC is supported by the European Union, Science Foundation of Ireland and Lero, the Science Foundation Ireland Research Centre for Software, grants #101028291 #13/RC/2094 and #SOWP2-TP0023, respectively. ME gratefully acknowledges the funding of a University of Warwick Chancellor’s Scholarship. EJB is grateful for the UKRI Future Leaders’ Fellowship funding (MR/V02261X/1) supporting this work. DOK is grateful for funding from Science Foundation Ireland (21/FFP-P/10160).
Appendix A ABAQUS formulation
@@
A.1 ABAQUS/Standard formulation
ABAQUS/Standard is an implicit solver where the model is updated from time $t$ to time $t+\Delta t$ using information from times $t$ and $t+\Delta t$. In this work, only static analyses are considered in ABAQUS/Standard. Therefore, since inertia is ignored, the relevant equations of motion are
@@ A.1 ABAQUS/Standard formulation
${\Delta\mathbf{u}_{\left(n+1\right)}}^{\left(t+\Delta t\right)}$ $=\left[{\mathbf{K}_{\left(n\right)}}^{\left(t+\Delta t\right)}\right]^{-1}\cdot\left({\mathbf{F}_{\left(n\right)}}^{\left(t+\Delta t\right)}\right),$ (3a)
${\mathbf{u}_{\left(n+1\right)}}^{\left(t+\Delta t\right)}$ $={{\mathbf{u}}_{(n)}}^{\left(t+\Delta t\right)}+\Delta{\mathbf{u}_{\left(n+1\right)}}^{\left(t+\Delta t\right)},$ (3b)
where the superscript $\left(t+\Delta t\right)$ is the time at which the equations are evaluated, the subscripts $\left(n\right)$ and $\left(n+1\right)$ indicate the current and next iterations respectively and $\mathbf{K}_{\left(n\right)}=\left(\partial\mathbf{F}/\partial\mathbf{u}\right)_{\left(n\right)}$ is the stiffness matrix at iteration $n$. The inversion of the stiffness matrix $\mathbf{K}_{\left(n\right)}$ is computationally expensive. The convergence of the Newton–Raphson method is measured by ensuring that all entries $\mathbf{F}(\mathbf{u})$ and $\Delta\mathbf{u}$ are sufficiently small [72]. At the end of time increment $t$, the next time increment $t+\Delta t$ is attempted, and load or displacement is incremented corresponding to $\Delta t$. If convergence is not achieved by ABAQUS in a certain time increment, $\Delta t$ is reduced and the Newton-Raphson method is attempted again. The reader should note that no velocities or accelerations are given by ABAQUS/Standard for static analyses. These can be manually extracted via postprocessing of the displacements (see B for more details).
A.2 ABAQUS/Explicit formulation
@@
$\dot{\mathbf{u}}^{\left(i+\frac{1}{2}\right)}$ $=\dot{\mathbf{u}}^{\left(i-\frac{1}{2}\right)}+\frac{\Delta t^{\left(i+1\right)}+\Delta t^{\left(i\right)}}{2}\ddot{\mathbf{u}}^{(i)},$ (5a)
${\mathbf{u}}^{\left(i+1\right)}$ $={\mathbf{u}}^{\left(i\right)}+\Delta t^{\left(i+1\right)}\dot{\mathbf{u}}^{\left(i+\frac{1}{2}\right)}.$ (5b)
@@
where $L^{e}$ is the smallest element dimension in the FE model, $\rho$ is the material density and $E$ is the Young’s modulus of the material. However, ABAQUS/Explicit enables the use of mass scaling. Mass scaling uniformly scales the density of the material within the simulation to facilitate longer time steps and reduce computational time [18]. The limited time increment means that explicit methods are conditionally stable, thus the mass scaling factor (MSF) must be chosen carefully to reduce simulation times without introducing inertial effects. The introduction of inertial effects can lead to erroneous and unstable outputs [73, 74, 62]. The general rule of thumb to assess inertial effects is to ensure that the ratio between the kinetic energy and the internal energy is small (5–10%) [67, 22].
Appendix B FE velocity calculation
ABAQUS/Standard does not output velocity components or PEEQ rate values for static analyses. Therefore, we manually calculate the velocity and PEEQ rate values. The velocity and PEEQ rate approximations are conducted via a Lagrangian time derivative of the displacements and the PEEQ values respectively. Since the procedure is identical for calculating both velocity and PEEQ rate values, we give details of the velocity calculation only.
@@ Appendix B FE velocity calculation
${v}^{(f)}_{x}$ $=\frac{\frac{1}{12}{u}^{(f-2)}_{x}-\frac{2}{3}{u}^{(f-1)}_{x}+\frac{2}{3}{u}^{(f+1)}_{x}-\frac{1}{12}{u}^{(f+2)}_{x}}{\Delta t},$ (7a)
${v}^{(f)}_{z}$ $=\frac{\frac{1}{12}{u}^{(f-2)}_{z}-\frac{2}{3}{u}^{(f-1)}_{z}+\frac{2}{3}{u}^{(f+1)}_{z}-\frac{1}{12}{u}^{(f+2)}_{z}}{\Delta t},$ (7b)
@@
where $\Delta t$ is constant between all time frames considered. In order to allow velocities and all other stress and strain components to be evaluated at the same time frame, all data in the main text is evaluated at a time point that is two time frames before $t=0.1\,$\mathrm{s}$$, which is the end of the rolling step. The authors ensured that $\Delta t$ is sufficiently small for accurate velocity approximations.
Appendix C Shear locking and incompatible CPE4I elements
Figure 14 shows shear stress evaluated at the integration points of CPE4 elements during a rolling simulation conducted as outlined in the main text with $N_{e}=40$ CPE4 elements through the half-thickness of the sheet. In Figure 14 a clear checkerboard pattern is observed, with more extreme shear stress values observed at the corners of elements. This is typical of shear locking behaviour [22].
@@ Appendix C Shear locking and incompatible CPE4I elements
Simulations with CPE4I and CPE4R elements employed do not exhibit such shear-locking behaviour. Figure 15 compares interpolated shear stress from two simulations: one with CPE4I elements and the other with CPE4R elements. Both simulations employ $N_{e}=15$ elements through the half-thickness, since simulations with $N_{e}\geq 20$ CPE4I elements required extremely small time increments which gave rise to convergence issues. The maximum relative error in Figure 15 is approximately 9% when compared to the maximum absolute shear stress in the roll gap. We expect that this error would reduce significantly as the mesh density increases. Given the expected similar shear stresses with increasing mesh density and that the simulation with CPE4I elements requires 17.7 CPU hours compared to 3.2 CPU hours for the simulation with CPE4R elements, CPE4I elements were not used for the main results in this work.
Figure 14: Shear stress evaluated at the integration points (i.e., no nodal averaging) for a rolling simulation using $N_{e}=40$ CPE4 elements through the half-thickness. There are significant jumps in shear stress across various elements, creating a checkerboard pattern, which is typical of shear-locking behaviour.
@@
Figure 15: Interpolated results for shear stress ($\mathrm{M}\mathrm{P}\mathrm{a}$) from simulations with CPE4I elements and CPE4R elements. Both simulations employ $N_{e}=15$ elements, a radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and $\varepsilon=0.125$. The results are almost identical and the CPU time is 17.7 hours for the CPE4I simulation compared to 3.2 hours for the CPE4R simulation. Readers should refer to Figure 3 which identifies the horizontal position for the solid, dotted and dashed lines, respectively.
Appendix D Relative error: choice of $x/L$ locations
@@
We detail here how the three $x/L$ positions used in the interpolated results throughout this paper were chosen. These $x/L$ positions represent the positions of maximum relative errors. To explain how these errors are calculated, let us consider the solid-line shear results in Figure 7(b). The interpolated shear result from the simulation with $N_{e}=5$ is subsequently interpolated onto a curve with 41 $z/h_{0}$ points. The absolute difference between this new curve and the interpolated shear result from the simulation with $N_{e}=40$ is calculated at each $z/h_{0}$ point. Finally, the absolute difference at each point is divided by the maximum absolute shear stress in the entire roll gap, as predicted by the simulation with $N_{e}=40$. At each $x/L$ location, the maximum relative error value out of the 41 values is stored in an array. The same procedure is carried out for $N_{e}=10$, $N_{e}=15$, $N_{e}=20$ and $N_{e}=30$, and all of these maximum errors are plotted in Figure 16.
@@
Figure 16: Relative error in (a) von Mises stress, (b) shear stress and (c) PEEQ (plastic equivalent strain) from simulations with varying mesh densities. All simulations are compared to a simulation with $N_{e}=40$ elements through the half-thickness. The error value at each $x/L$ location represents the maximum error through the thickness of the sheet at that $x/L$ position. All simulations employ a radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and $\varepsilon=0.125$.
@@
Figure 16 shows that $x/L=0.025$ is a position where high relative errors occur, and hence this is one of the $x/L$ positions used for interpolations in this paper. Figure 16(a) and Figure 16(b) also show a spike in relative errors near $x/L=1.09$. This is another $x/L$ position used to calculate the interpolated results. It is not clear from Figure 16(b) and Figure 16(c) which third $x/L$ position to choose. However, Figure 16(a) shows a small peak in relative error around the neutral point at $x/L=0.6913$ (the neutral point value of $x/L=0.6913$ is predicted by the implicit simulation with $N_{e}=40$). Given this peak and the interesting dynamics around the neutral point in general, $x/L=0.6913$ is the third position used in the interpolated results throughout this paper.
Appendix E Slip line theory
@@
$=\begin{pmatrix}S_{xx}-p&S_{xz}\\ S_{xz}&S_{zz}-p\end{pmatrix},$
@@
$0=\frac{\partial\sigma_{xx}}{\partial x}+\frac{\partial\sigma_{xz}}{\partial z}$ $=-\frac{\partial p}{\partial x}-\frac{2Y}{\sqrt{3}}\!\left(\!\cos{2\theta}\frac{\partial\theta}{\partial x}+\sin{2\theta}\frac{\partial\theta}{\partial z}\right)\!,$ (12a)
$0=\frac{\partial\sigma_{xz}}{\partial x}+\frac{\partial\sigma_{zz}}{\partial z}$ $=-\frac{\partial p}{\partial z}-\frac{2Y}{\sqrt{3}}\!\left(\!\sin{2\theta}\frac{\partial\theta}{\partial x}-\cos{2\theta}\frac{\partial\theta}{\partial z}\right)\!.$ (12b)
Equations 12a and 12b are hyperbolic PDEs and can be solved by the method of characteristics, (see Johnson et al. [36] for a detailed description). The characteristic directions of each of these equations are given by
$\dfrac{dz}{dx}$ $=\tan{\theta}$ and $\dfrac{dz}{dx}$ $=\tan{\left(\theta+\frac{\pi}{2}\right)}$ (13)